[figure: 9.7 cm]

MANIFESTATION OF SPIN-GLASS-LIKE BEHAVIOR IN THE

ORGANOMETALLIC MAGNET

V.P. Dyakonov^{1,3}, M. Arciszewska^{1}, W. Dobrowolski^{1}, A. Szewczyk^{1}, H. Szymczak^{1}, E. Aksymentieva^{2}, V. Shapovalov^{3}, E. Zubov^{3} 1Institute of Physics, Warszawa, Poland 2 Lviv State University, Ukraine 3 Donetsk Physico-Technical Institute, Ukraine

Ac susceptibility and dc magnetization of the organometallic magnet Na[FeO_{6}(C_{10}H_{8}N)_{3}] have been studied in temperature range of 4.2–240 K, in external magnetic field up to 90 kOe. Some peculiarities of magnetic behavior which are characteristic of spin-glasses, e.g. cusps in the susceptibility and ZFC magnetization appearing at the freezing temperature, a field–dependent irreversibility between the FC and ZFC magnetization, and dependencies of the T_{cusp} temperature on the frequency and magnetic field strength, were observed at low temperatures.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Vladimir P.Dyakonov

Address for correspondence : Al.Lotników 32/46 02-668,Institute of Physics, Warszawa, Poland

Email address : dyakon@ifpan.edu.pl